OpenStudy (anonymous):

tan(52 degrees)=15/y solve for y

OpenStudy (anonymous):

\[a=\frac{b}{y}\] solve for \(y\)

myininaya (myininaya):

I would reciprocal both sides first yes I used that one word as a verb

OpenStudy (anonymous):

\[15=\frac{2}{y}\] solve for \(y\)

OpenStudy (anonymous):

reciprocal both sides??

myininaya (myininaya):

assume a and b don't equal zero and assuming a=b then 1/a=1/b

myininaya (myininaya):

but yeah if you can solve a=b/y for y then that is a good start

OpenStudy (anonymous):

so would I multiply the y on both sides so 15*y=2?

myininaya (myininaya):

yeah

myininaya (myininaya):

so here you would multiply y on both sides as well y tan(52)=15

OpenStudy (anonymous):

ok thanks a bunch!